OpenStudy (pottersheep):

Please help. Find the dimensions of a rectangle that has a perimeter of 16cm and area of 13 cm^2

OpenStudy (pottersheep):

still here ~

OpenStudy (anonymous):

1) P = 2b + 2h = 16 2) A = b*h = 13 From eq. (1): b = 8 - h isolate in eq. (2) : (8 - h)(h) = 13 8h - h^2 = 13 h^2 - 8h + 13 = 0 using quadratic formula: \[(-b +/- \sqrt{b^2 - 4ac}) / 2a\] h = 4 +/- sqrt(3) so assume h = 4+ sqrt(3) using eq. (1): b = 8 - (4+sqrt(3)) b = 4 - sqrt(3) this can be verified in eq. (2) also, the assumption of h = 4 + sqrt(3) instead of 4 - sqrt(3) was arbitrary, the dimensions will be the same independent of this assumption so the dimentions are 4 + sqrt(3), 4 - sqrt(3)